If T is nilpotent transformation on V of index m. Prove that m cannot exceed the dimension of V


1
Expert's answer
2021-07-07T16:05:27-0400

Question

If T is nilpotent transformation on V of index m. Prove that m cannot exceed the dimension of V.

Solution;

If T is a linear transformation on V of order m, then dim V "\ge" m by induction.

Assume the claim herein to be true for index m-1, and let W={x"\epsilon" V|Tm-1x=0}

W is the null space of Tm-1

Clearly,T maps W onto itself and T is nilpotent of order m-1.

By induction W"\ge" m-1 since W is a proper subspace of V.

We then conclude that ,the minimal polynomial of T is xm and therefore

m"\le" dim V